OpenStudy (kikuo):

Can someone explain why we flip fractions when we divide fractions? I know how to and that multiplication is the opposite of division, but I just can't seem to make sense of it no matter where I look.

OpenStudy (jdoe0001):

\(\large \cfrac{\frac{a}{b}}{\frac{c}{{\color{blue}{ d}}}}\implies \cfrac{a}{b}\cdot \cfrac{{\color{blue}{ d}}}{c} ?\) so...what are you asking again?

OpenStudy (caozeyuan):

what is 5/5? and what is 5*1/5?

OpenStudy (caozeyuan):

so if you have to calculate a/b, you can do this\[\frac{ a }{ b }=\frac{ a*1 }{ b }=a*\frac{ 1 }{b }\]

OpenStudy (kikuo):

@jdoe0001 In words, can you describe what you just posted? @caozeyuan Same for you. :)

OpenStudy (caozeyuan):

if a divides b, we can say a times 1 then divides b, since a times 1 is a

OpenStudy (caozeyuan):

and it is the same as a times (1 divides b), which is a time the inverse of b

OpenStudy (caozeyuan):

one of the skills you need to acquire is to think aabstractly, i.e. do not rely on words but on equations, since equation is less confusing

OpenStudy (kikuo):

I don't understand the last step. @caozeyuan Also, how does this explain why we flip one fraction?

OpenStudy (mathstudent55):

For example, dividing by 2 is the same as multiplying by 1/2. Since when you flip 2 (written as 2/1) you get 1/2, to divide by 2, you flip 2 into 1/2 and you multiply by1/2.

OpenStudy (mathstudent55):

Start with a number a, and you want to divide by b, .\(b \ne 0\): \(a \div b\) A fraction means a division, so \(a \div b\) is the same as \(\dfrac{a}{b} \) Let's rewrite our division as a fraction: \(a \div b = \dfrac{a}{b} \) From the multiplication of fractions, we know that \(\dfrac{a}{b} = \dfrac{a}{1} \times \dfrac{1}{b} \) Our original division, \(a \div b\) can also be written as \(\dfrac{a}{1} \div \dfrac{b}{1} \) Now we can conclude that \(\dfrac{a}{1} \div \dfrac{b}{1} = \dfrac{a}{1} \times \dfrac{1}{b} \)

OpenStudy (kikuo):

@mathstudent55 Your route makes sense to me the most. I'm confused on a few things, but I'm starting to understand now. For example, dividing by 2 is the same as multiplying by 1/2. Since when you flip 2 (written as 2/1) you get 1/2, to divide by 2, you flip 2 into 1/2 and you multiply by1/2. From the multiplication of fractions, we know that These two parts confused me.

OpenStudy (skullpatrol):

What grade level are you?

OpenStudy (kikuo):

I'm in 10th. I know how to divide fractions, but I want to understand why we divide fractions the way we do. Our teacher never taught us. We were simply taught when you divide fractions you flip the second fraction and multiply, but he didn't explain why we did this.

OpenStudy (kikuo):

So, out of curiosity, I'm trying to understand it.

OpenStudy (skullpatrol):

Ok, give me a minute :-)

OpenStudy (kikuo):

Oh thanks! Haha, I was worried someone wouldn't continue.

OpenStudy (mathstudent55):

I think it can be explained simply as @caozeuyan did way above. \(a \div b = \dfrac{a}{b} = \dfrac{a \times 1}{b} = a \times \dfrac{1}{b} \)

OpenStudy (skullpatrol):

We "flip" fractions because of the definition of division: $$\Huge a \div b = \dfrac{a}{b} = a \cdot \dfrac{1}{b}$$ This means we must multiply by the reciprocal of b. $$\dfrac{1}{\dfrac{a}{b}} = \dfrac{b}{a}$$ Proof: a/b • b/a = a • 1/b • b • 1/a by the definition of division = a • 1/a • b • 1/b by the comm and assoc prop of mult = 1 • 1 by the property of reciprocals = 1 by the identity property of mult But a/b • 1/ (a/b) = 1 by the property of reciprocals Thus a/b • 1/ (a/b) = a/b • b/a substitution principle from above 1/ (a/b) = b/a by the division property of equality

OpenStudy (kikuo):

@skullpatrol This means we must multiply by the reciprocal of b. I'm stuck on this part.

OpenStudy (skullpatrol):

You must multiply by the the reciprocal of the denominator, right?

OpenStudy (kikuo):

I'm extremely confused. http://prntscr.com/ahvjqs I'm not sure where you're getting this form from haha. I see where the 1/a comes in but I don't know how that =b/a and I don't know what the other b there is for.

OpenStudy (skullpatrol):

The a and b in the large part at the top is just the definition, that we all memorize. Ok?

OpenStudy (kikuo):

Alright, so what about this part? http://prntscr.com/ahvn95

OpenStudy (skullpatrol):

I am considering the large b in the definition as a/b.

OpenStudy (skullpatrol):

I should have used different letters, sorry

OpenStudy (kikuo):

http://prntscr.com/ahvo3z This one? LOL why don't we just start over?

OpenStudy (skullpatrol):

Your question was "why do we flip the fraction" when we divide, right?

OpenStudy (kikuo):

Yep : )

OpenStudy (skullpatrol):

And you know that to divide by a number you multiply by its reciprocal, right?

OpenStudy (kikuo):

Yes

OpenStudy (skullpatrol):

So let's say we need to divide 1 by (a/b).

OpenStudy (skullpatrol):

For example 1/(2/3)

OpenStudy (kikuo):

Why can we do that? 1/2/3

OpenStudy (skullpatrol):

Do what?

OpenStudy (skullpatrol):

What does 1 / ( 2/3 ) = ?

OpenStudy (skullpatrol):

$$\Huge \dfrac{1}{\dfrac{2}{3}}$$

OpenStudy (skullpatrol):

Put the numbers in your calculator and find the answer please.

OpenStudy (kikuo):

Oh, we can't solve by hand? @skullpatrol

OpenStudy (skullpatrol):

Sure.

OpenStudy (skullpatrol):

What do you get?

OpenStudy (kikuo):

1/1 divided by 2/3

OpenStudy (skullpatrol):

Yes

OpenStudy (skullpatrol):

Sure.

OpenStudy (skullpatrol):

Put the numbers in your calculator and find the answer please.

OpenStudy (kikuo):

0.16

OpenStudy (skullpatrol):

You should get 1.5 = 3/2

OpenStudy (anonymous):

are you trying to do it or to make sense out of it?

OpenStudy (kikuo):

I'm trying to make sense out of it, and I'm extremely confused.

OpenStudy (kikuo):

That's weird, I got 0.16 again.

OpenStudy (skullpatrol):

Divide 2/3

OpenStudy (skullpatrol):

2 divided by 3 = ?

OpenStudy (kikuo):

Got 0.6

OpenStudy (skullpatrol):

0.666...

OpenStudy (kikuo):

Ah, I didn't know I was supposed to reveal the whole thing.

OpenStudy (skullpatrol):

Now take 1 and divide it by 0.666...

OpenStudy (anonymous):

can you think of an example of dividing by a fraction? here is a simple one how many quarters are in two dollars?

OpenStudy (kikuo):

4 quarters make one dollar so I assume 8 would make 2

OpenStudy (anonymous):

yes what did you do? you multiplied 2 by 4

OpenStudy (kikuo):

Got 0.666 Yep, that's what I did.

OpenStudy (anonymous):

so as you see \[2\div \frac{1}{4}=2\times 4=8\]

OpenStudy (anonymous):

want another example?

OpenStudy (mathstudent55):